2. Verify that the point P(a cos θ, b sin θ) lies on the ellipse
x^2/a^2 + y^2/b^2 = 1,
where a and b are the semi-major and semi-minor axes respectively of the ellipse .
Find thegradient of the tangent to the curve at P and show that the equation of the normal at P is
ax sin θ − by cos θ = (a^2 − b^2) sin θ cos θ.

If P is not on the axes and if the normal at P passes through the point B(0, b), Show that a^2 > 2b^2.

If further, the tangent at P meets the y-axis at Q, show that
|BQ| =a^2/b^2 .
1
Expert's answer
2020-06-30T17:24:13-0400

equation of ellipse x2/a2 + y2/b2 = 1

we have points P(acos"\theta" , bsin"\theta" )

put the points in the equation of ellipse

take L.H.S

(acos"\theta" )2/a2 + (bsin"\theta" )2/b2

a2cos2"\theta" /a2 + b2sin2"\theta"/b2

cos2"\theta" + sin2"\theta"

=1

hence these point are lies on the ellipse


the equation of normal is at p is ax sin"\theta" − by cos"\theta" = (a2 − b2) sin"\theta" cos"\theta"

the gradient of the equation is dy/dx

so the gradient is

a sin"\theta" {dy/dx(x)} - by cos"\theta" {dy/dx(1)} =(a2-b2)sin"\theta" cos"\theta" {dy/dx(1)}

asin"\theta" - 0 = 0

asin"\theta" = 0 is the gradient of the tangent to the curve at P

 


equation of normal is

ax sin"\theta" - by cos"\theta" = (a2-b2)sin"\theta" cos"\theta"

at point B(0,b)

a (0) sin"\theta" - b (b)cos"\theta" =(a2-b2)sin"\theta" cos"\theta"

0-b2cos"\theta" =(a2-b2)sin"\theta" cos"\theta"

hence a2 > 2b2


if the tangent P meets the y axis at Q

than |BQ| = a2/b2

the point B (0,b) similarly at y axis the point Q (a,0)

so the the point BQ(a,b)

hence |BQ| = a2/b2